# Helping or Herding? Reward Model Ensembles Mitigate but do not Eliminate Reward Hacking

### v1

@@ 5 When do Reward Model Ensembles Fail?
To demonstrate this, we manually analyze ensemble outputs to detect frequent errors, and then perform a qualitative analysis. Figure˜9 shows the results of this analysis on all three benchmarks. The x-axis corresponds to outputs of the model after training for a certain number of steps, and the y-axis is a statistic of interest (e.g., average output length). We plot the statistic value for the pretrained ensemble (using mean as a representative aggregation function) and for its members. In addition, for tl;dr and helpfulness, where the reward model is trained on the preference data, we show the statistic value on the preference data validation set, conditioned on the label ‘Preferred’ or ‘Rejected’.
  • For helpfulness (Figure˜9(a)), outputs tend to be in a format of a list, and thus we write a regular expression that captures this format. The fraction of outputs that have this pattern increases to roughly 50% for 3 members of the ensemble and to the ensemble itself. Looking at the preference data, we do not detect a tendency to produce list outputs in the preferred responses, as the fraction of outputs that matches this format is roughly 8% for both the preferred and rejected responses.
  • For tl;dr (Figure˜9(b)), RLHF alignment leads to longer summaries (Singhal et al., 2023) and also outputs that are more extractive, i.e., copy more from the input. Summary length in characters grows substantially for the ensemble and all its members, where for the ensemble, length increases by a factor of two. On the preference data, indeed preferred responses are slightly longer than rejected responses, but much shorter than outputs post-RLHF. We also compute the longest common subsequence (in characters) between the document and the summary and find that it increases for the ensemble from 28.2 to 49.1. Again, the tendency for copying from the document already occurs in the preference data to a small degree, but is amplified by RLHF. [^3] The distribution of outputs in the preference data is not identical to the distribution of outputs before RLHF, and therefore the statistics after zero training steps do not necessarily match those of the preference data.
  • For xsum/nli (Figure˜9(c)), training for factuality tends to make summaries shorter. Additionally, precise numbers are typically omitted from the summaries. Figure˜9 shows how all members of the ensemble and the ensemble itself exhibit this phenomenon, with length in characters decreasing rapidly, as well as the fraction of examples that contain any numeric value whatsoever.
@@ Appendix A Numerical results for best-of-n reranking
Average agreement between reward models are shown in Tables 3-6. Autoevaluation results are shown in Table˜7 and 8.
@@
Table 3: tl;dr best-of-n agreement
@@
Table 4: helpfulness best-of-n agreement
@@
Table 6: helpfulness top 1 agreement
@@
base finetune mean 0.220-0.220 $0.700$
mean minus std 0.186-0.186 $0.703$
median 0.231-0.231 $0.700$
min 0.177-0.177 $0.710$
pretrain mean 0.130-0.130 $0.721$
mean minus std 0.086-0.086 $0.731$
median 0.155-0.155 $0.715$
min 0.086-0.086 $0.727$
single RM single RM 0.244-0.244 $0.685$
@@ Appendix B Numerical results for RLHF
Full RLHF numerical results for helpfulness and tl;dr are shown in Table˜9 and Table˜10.
@@
Table 9: helpfulness RLHF numerical results
@@
Table 10: tl;dr RLHF numerical results
@@ Appendix C Hyperparameters
We provide the hyperparameters for reward model and RLHF training in Table˜11 and Table˜12. For reward models, we use the validation set to choose the best checkpoint along training. For RLHF, we take the last checkpoint.
@@
Table 11: Hyper-parameters for reward model training
@@
Table 12: Hyper-parameters for RLHF

### v2

@@ 5 When do Reward Model Ensembles Fail?
To demonstrate this, we manually analyze ensemble outputs to detect frequent errors, and then perform a qualitative analysis. Figure 9 shows the results of this analysis on all three benchmarks. The x-axis corresponds to outputs of the model after training for a certain number of steps, and the y-axis is a statistic of interest (e.g., average output length). We plot the statistic value for the pretrained ensemble (using mean as a representative aggregation function) and for its members. In addition, for tl;dr and helpfulness, where the reward model is trained on the preference data, we show the statistic value on the preference data validation set, conditioned on the label ‘Preferred’ or ‘Rejected’.
  • For helpfulness (Figure 8(a)), outputs tend to be in a format of a list, and thus we write a regular expression that captures this format. The fraction of outputs that have this pattern increases to roughly 50% for 3 members of the ensemble and to the ensemble itself. Looking at the preference data, we do not detect a tendency to produce list outputs in the preferred responses, as the fraction of outputs that matches this format is roughly 8% for both the preferred and rejected responses.
  • For tl;dr (Figure 8(b)), RLHF alignment leads to longer summaries (Singhal et al., 2023) and also outputs that are more extractive, i.e., copy more from the input. Summary length in characters grows substantially for the ensemble and all its members, where for the ensemble, length increases by a factor of two. On the preference data, indeed preferred responses are slightly longer than rejected responses, but much shorter than outputs post-RLHF. We also compute the longest common subsequence (in characters) between the document and the summary and find that it increases for the ensemble from 28.2 to 49.1. Again, the tendency for copying from the document already occurs in the preference data to a small degree, but is amplified by RLHF. [^2] The distribution of outputs in the preference data is not identical to the distribution of outputs before RLHF, and therefore the statistics after zero training steps do not necessarily match those of the preference data.
  • For xsum/nli (Figure 8(c)), training for factuality tends to make summaries shorter. Additionally, precise numbers are typically omitted from the summaries. Figure 9 shows how all members of the ensemble and the ensemble itself exhibit this phenomenon, with length in characters decreasing rapidly, as well as the fraction of examples that contain any numeric value whatsoever.
@@ Appendix A Numerical results for best-of-n reranking
Average agreement between reward models are shown in Tables 3-6. Autoevaluation results are shown in Table 7 and 8.
@@
Table 3: tl;dr best-of-n agreement.
@@
Table 4: helpfulness best-of-n agreement.
@@
Table 6: helpfulness top-1 agreement.
@@
base finetune mean $-0.220$ $0.700$
mean minus std $-0.186$ $0.703$
median $-0.231$ $0.700$
min $-0.177$ $0.710$
pretrain mean $-0.130$ $0.721$
mean minus std $-0.086$ $0.731$
median $-0.155$ $0.715$
min $-0.086$ $0.727$
single RM single RM $-0.244$ $0.685$
@@ Appendix B Numerical results for RLHF
Full RLHF numerical results for helpfulness and tl;dr are shown in Table 9 and Table 10.
@@
Table 9: helpfulness RLHF numerical results.
@@
Table 10: tl;dr RLHF numerical results.
@@ Appendix C Hyperparameters
We provide the hyperparameters for reward model and RLHF training in Table 11 and Table 12. For reward models, we use the validation set to choose the best checkpoint along training. For RLHF, we take the last checkpoint.
@@
Table 11: Hyper-parameters for reward model training.
@@
Table 12: Hyper-parameters for RLHF.